OpenStudy (anonymous):

Write the equation of a line that passes through the point (-2, 1) and has a slope of 4. Select one of the options below as your answer: A. y = 4x − 7 B. y = 4x + 1 C. y = 4x + 3 D. y = 4x + 9

OpenStudy (anonymous):

@hartnn

OpenStudy (aroub):

y=mx+b; x=-2, y=1, m=4 1=4(-2)+b 1=-8+b 9=b => y=4x+9